OpenStudy (anonymous):

Solve the system 3x – 2y = 7 and 9x – 6y = 3 by using graph paper or graphing technology. What is the solution to the system?

OpenStudy (anonymous):

It doesn't have a solution 3(3x - 2y) = 3(7) => 9x - 6y = 21... Contradicts 9x - 6y = 3. If you graph 9x - 6y and 3x - 2y, your lines will be parrallel. To graph it and see for yourself. Solve for y: -2y = 3 - 3x ... y = -(3/2) - -(3/2)x = (3/2)x - 3/2 or -6y = 21 - 9x ... y = -(7/2) - -(9/6)x = (3/2)x - 7/2 Notice they both have the same slope. Only the intercept differs.